# Simulation of Topological $X$-Gates via Braiding of Majorana Zero Modes in an Interacting Quantum Dot System

### v1

@@ Abstract
Recent advances in quantum dot platforms have opened new pathways for realizing Majorana zero modes (MZMs) and simulating topological quantum computation. Here we propose an experimentally feasible setup for implementing topological $\sqrt{X}$- and $X$-quantum gates in an interacting $Y$-shaped quantum-dot array. The proposed novel architecture enables both braiding and charge readout through simple fusion operations controlled by gate-tunable potentials. Using many-body time-dependent simulations based on exact diagonalization, we analyze the braiding and fusion dynamics of MZMs in the presence of nearest-neighbor Coulomb interactions and pairing disorder. We compute diabatic errors, braiding fidelity, and the time- and space-resolved electron and hole components of the local density of states to monitor the braiding process. Our results show that even weak interactions or pairing disorder induce oscillations in the braiding fidelity, thereby setting an upper bound on the braiding speed. Furthermore, we demonstrate that comparing fusion outcomes before and after braiding provides a direct and experimentally accessible signature of the non-Abelian nature of MZMs in quantum dot systems.
@@ Introduction
Quantum-dot arrays offer a promising platform for simulating topological quantum computation and testing non-Abelian statistics through controlled exchange (braiding) of MZMs Boross2024 ; Tsintzis2024 ; Miles2025 . In quantum-dot systems, Majorana zero modes (MZMs) are localized on-site at the sweet spot Dvir ; Bordin , enabling fusion and braiding even in comparatively small system sizes, in contrast to conventional nanowire-based platforms. The experimental realization of MZMs braiding can provide definitive evidence of non-Abelian statistics and, importantly, distinguish them from unwanted trivial Andreev bound states. Probing such non-Abelian properties requires the adiabatic exchange of MZMs using minimal geometries such as $T$- or $Y$-shaped quantum-dot networks Boross2024 ; Tsintzis2024 ; Pandey2025 ; Pandey2023 . The controlled motion of MZMs has been demonstrated experimentally in three-site quantum dot configurations tenHaaf , and numerically in six-dot systems Pandey2025_PRB . Realistic modeling of such braiding protocols demands full many-body, time-dependent simulations that incorporate the effects of both the degenerate ground states and bulk excitations. While prior studies on proximitized nanowire systems often relied on non-interacting Harper ; Amorim2015 or low-energy effective models, braiding in next-generation quantum-dot architectures remains less explored.
@@
Figure 1: Initialization, braiding, and readout of a Majorana zero-mode qubit in a $Y$-shaped quantum-dot array. Schematic of a device comprising seven quantum dots coupled via semiconductor–superconductor hybrid structures arranged in a $Y$-shaped geometry. The red dots indicates the non-topological regime ($\mu_{j}=\mu_{0}$), while the blue dots represent topological sites ($\mu_{j}=0$). a, At time $t=0$, two pairs of Majorana zero modes (MZMs), $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$, are localized at the ends of the second and third legs, respectively. b, $Z$-fusion operations, $F^{Z}_{ba}$ and $F^{Z}_{cd}$, are performed by tuning the on-site chemical potentials $\mu_{j}(t)$ at sites $j=4$ and $6$. The fusion outcomes are detected through charge sensing by nearby quantum-dot sensors (yellow) coupled to sites $j=5$ and $7$. c, Illustration of the double braiding operation $\hat{U}_{bc}^{2}$, exchanging MZMs $\gamma_{b}$ and $\gamma_{c}$. d, Final readout is performed via $Z$-fusion of $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$ after the double braiding process.
In this work, we propose a novel scheme to test non-Abelian statistics in an interacting quantum-dot system arranged in a $Y$-shaped geometry (Fig. 1). Our experimentally feasible setup, composed of only a few quantum dots, enables the initialization, braiding, and readout of a MZM-based qubit within its coherence lifetime Bordin . We perform dynamical simulations of an $X$-gate in an interacting quantum-dot system using the full many-body wavefunction. We simulate the braiding of MZMs in interacting quantum dots arranged in a $Y$-shaped geometry by solving the time-dependent many-body Schrödinger equation via exact diagonalization. To assess the fidelity of the braiding protocol, we compute the diabatic error and transition probabilities. We visualize the positions of the MZMs during the $X$-gate process by calculating the time-, space-, and energy-resolved local density of states, $LDOS(t,j,\omega)$ Pandey2023_PRB , which should be accessible in time-dependent differential conductance measurements using scanning tunneling spectroscopy Houselt2010 .
@@ Results
Here, we introduce a minimal and experimentally feasible setup consisting of seven interacting quantum dots arranged in a $Y$-shaped geometry (see Fig. 1), suitable for implementing MZM braiding operations and readout. In this configuration, the quantum dots are connected via electrostatically tunable gates, which enable dynamic control over the hopping amplitude $t^{h}(t)$, pairing strength $\Delta(t)$, and on-site chemical potential $\mu_{j}(t)$ Dvir ; Bordin ; Liu2025 ; Pandey2024_PRR , each carrying a time dependence. This high degree of tunability allows for the adiabatic transport of MZMs from one quantum dot to another tenHaaf ; Pandey2025_PRB , thereby enabling both braiding and fusion operations of MZMs. The Hamiltonian for the $Y$-shaped quantum dot array (with $\phi_{I}=0$, $\phi_{II}=\pi/6$, and $\phi_{III}=-\pi/6$ for each leg Harper ) can be divided into four distinct parts. The Hamiltonian for each arm is given by:
@@
Here, l=(L1)/3l=(L-1)/3, where $L$ is the total number of sites, $c_{j}$ is the fermionic annihilation operator, and $n_{j}=c^{\dagger}_{j}c_{j}$ is the number operator. The parameters $t^{h}_{ij}$ and $\Delta_{ij}(t)$ denote the bond-dependent hopping and pairing amplitudes, respectively. We assume $t^{h}=\Delta$ for our numerical calculations (unless stated otherwise). We use a smooth, time-dependent sine-squared ramp protocol Pandey2025_PRB , given by
@@
Figure 3: Simulation of Majorana braiding and qubit dynamics in a $Y$-shaped quantum-dot array. a, Initial configuration at $t=0$ with two pairs of spatially separated Majorana zero modes (MZMs), $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$. b–g, Space- and time-resolved local density of states $LDOS(\omega,t,j)$ at different stages of the braiding sequence. b, At $t=0$, four zero-energy peaks are observed on sites $j=4,5,6,7$, corresponding to the four MZMs. c, At $t=2T/6$, the MZM $\gamma_{b}$ is transferred from site $j=4$ to $j=1$, leading to zero-energy peaks at $j=1,5,6,7$. d, At $t=3T/6$, $\gamma_{c}$ becomes partially delocalized, resulting in finite spectral weight at sites $j=3,4,6$, while sharp peaks remain at $j=1,5,6$. e, At $t=4T/6$, $\gamma_{c}$ becomes localized at site $j=4$, and zero-energy peaks are observed at $j=1,4,5,7$. f, At $t=T$, after completing a single braid, $\gamma_{b}$ and $\gamma_{c}$ are exchanged, as indicated by the restored peaks at $j=4,5,6,7$. g, At $t=2T$, following a full double braid, the system returns to its initial configuration, showing four zero-energy peaks identical to the initial state. h, Time evolution of the diabatic error $f(t)=1-\sum_{n=1}^{4}|\langle\psi(t)|\psi_{n}^{I}\rangle|^{2}$ as a function of scaled time $t/\tau$ for different ramp durations $\tau$. i, Time evolution of the transition probabilities $P_{n}(t)=|\langle\psi_{n}(0)|\psi(t)\rangle|^{2}$ for all four eigenstates ($n=1,2,3,4$), shown for $\tau=50$. j, Final transition probability $P(2T)$ at the end of the protocol ($t=t_{f}$) as a function of switching time $\tau$ for different values of the Coulomb interaction strength $V$.
@@
where $\psi_{n}^{I}$ denotes the instantaneous many-body eigenstates of the system. For small values of $\tau$ (corresponding to fast switching), the diabatic error takes a larger value, indicating transitions to higher-energy states above the gap. Interestingly, we find that for $\tau\geq 40$, the diabatic error remains low ($f(t)\sim 10^{-4}$), showing neglegible excitation above the ground-state manifold. For larger $\tau$, the spikes in the diabatic error appear at times when the on-site potentials at the three initial sites $(j=1,2,3)$ are changed simultaneously. The diabatic error for smaller values of switching time leads to larger braiding error and gives a lower bound to the switching time $\tau\gtrsim\hbar/\Delta_{gap}$ Knapp2016 ; Peeters2024 (here $\Delta_{gap}$ is the excitation gap in the topological phase).
@@
Next, we perform a double braiding operation $U^{2}_{bc}$ between MZMs $\gamma_{b}$ and $\gamma_{c}$ (see schematic of Fig. 4(b)). As discussed earlier, this operation transforms the MZM qubit state from $|11\rangle$ to $|00\rangle$. To read out the final state after double braiding (at $t=2T$), we repeat the same $Z$-fusions $F^{Z}_{ba}$ and $F^{Z}_{cd}$. The second row of Figs. 4(g-j) presents the results of $Z$-fusion for different fusion switching times $\tau_{f}$, following a braiding protocol with $\tau=50$ (braiding switching time). In this case, the hole component of the local density of states, $LDOS^{h}(\omega,t,j)$, exhibits a peak with height one at $\omega=0$ (Fig. 4(h)), while the electron component, $LDOS^{e}(\omega,t,j)$, shows no peak at site $j=5$ (Fig. 4(g)). This indicates the formation of a vacuum state ($|00\rangle$) after $Z$-fusion, in contrast to the electron formation $|11\rangle$ observed prior to braiding. Figures 4(i,j), show the time-dependent charge density $\langle n_{j}(t)\rangle$ at sites $j=5$ and $j=7$, respectively, for various $\tau_{f}$ (fusion switching time). In both cases, the charge density approaches zero for $\tau_{f}\gtrsim 40$, confirming the formation of holes at both sites following fusion. These results validate that the double braid operation leads to a state transformation $|11\rangle\rightarrow|00\rangle$ and should be obtained in the quantum-dot experiment using charge sensor Szechenyi2020 ; Gharavi2016 .
@@ Discussion
Acknowledgments
@@
bradraj.pandey@gmail.com.

### v2

@@ Abstract
Recent advances in quantum dot platforms have opened new platform for realizing Majorana zero modes (MZMs) and simulating topological quantum computation. Here we propose an experimentally feasible setup for implementing topological $\sqrt{X}$- and $X$-quantum gates in an interacting $Y$-shaped quantum-dot array. The proposed novel architecture enables both braiding and charge readout through simple fusion operations controlled by gate-tunable potentials. Using many-body time-dependent simulations based on exact diagonalization, we analyze the braiding and fusion dynamics of MZMs in the presence of nearest-neighbor Coulomb interactions and pairing disorder. We compute diabatic errors, braiding fidelity, and the time- and space-resolved electron and hole components of the local density of states to monitor the braiding process. Our results show that even weak interactions or pairing disorder induce oscillations in the braiding fidelity, thereby setting an upper bound on the braiding speed. Furthermore, we demonstrate that comparing fusion outcomes before and after braiding provides a direct and experimentally accessible signature of the non-Abelian nature of MZMs in quantum dot systems.
@@ Introduction
Quantum-dot arrays offer a promising platform for simulating topological quantum computation and testing non-Abelian statistics through controlled exchange (braiding) of MZMs Boross2024 ; Tsintzis2024 ; Miles2025 . In quantum-dot systems, Majorana zero modes (MZMs) are localized on-site at the sweet spot Dvir ; Bordin , enabling fusion and braiding even in comparatively small system sizes, in contrast to conventional nanowire-based platforms. The experimental realization of MZMs braiding can provide definitive evidence of non-Abelian statistics and, importantly, distinguish them from unwanted trivial Andreev bound states. Probing such non-Abelian properties requires the adiabatic exchange of MZMs using minimal geometries such as $T$- or $Y$-shaped quantum-dot networks Boross2024 ; Tsintzis2024 ; Pandey2025 ; Pandey2023 ; Sekania2017 . The controlled motion of MZMs has been demonstrated experimentally in three-site quantum dot configurations tenHaaf , and numerically in six-dot systems Pandey2025_PRB . Realistic modeling of such braiding protocols demands full many-body, time-dependent simulations that incorporate the effects of both the degenerate ground states and bulk excitations. While prior studies on proximitized nanowire systems often relied on non-interacting Harper ; Amorim2015 or low-energy effective models, braiding of four MZMs in next-generation quantum-dot architectures remains less explored Sekania2017 ; Miles2025 .
@@
Figure 1: Initialization, braiding, and readout of a Majorana zero-mode qubit in a $Y$-shaped quantum-dot array. Schematic of a device comprising seven quantum dots coupled via semiconductor–superconductor hybrid structures arranged in a $Y$-shaped geometry. Red dots indicate non-topological sites ($\mu_{j}=\mu_{0}$), while blue dots denote topological sites ($\mu_{j}=0$). a, At $t=0$, two pairs of Majorana zero modes (MZMs), $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$, are localized at the ends of the second and third legs, respectively. b, $Z$-fusion operations, $F^{Z}_{ba}$ and $F^{Z}_{cd}$, are performed by tuning the on-site chemical potentials $\mu_{j}(t)$ at sites $j=4$ and $6$. The fusion outcome is an electron $\psi$ on sites $j=5$ and $6$. The outcomes are detected through charge sensing by nearby quantum-dot sensors (yellow) coupled to sites $j=5$ and $7$. c, Illustration of the double-braiding operation $\hat{U}_{bc}^{2}$, exchanging MZMs $\gamma_{b}$ and $\gamma_{c}$. d, Final readout is performed via $Z$-fusion of $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$ after the double-braiding process. The final fusion outcome is the vacuum II on sites $j=5$ and $6$, which differs from the initial fusion outcomes.
In this work, we propose a novel scheme to test non-Abelian statistics in an interacting quantum-dot system arranged in a $Y$-shaped geometry (Fig. 1). Our experimentally feasible setup, composed of only a few quantum dots, enables the initialization, braiding, and readout of a MZM-based qubit within its coherence lifetime Bordin . We perform dynamical simulations of an $X$-gate in an interacting quantum-dot system using the full many-body wavefunction. We simulate the braiding of MZMs in interacting quantum dots arranged in a $Y$-shaped geometry by solving the time-dependent many-body Schrödinger equation via exact diagonalization. To assess the fidelity of the braiding protocol, we compute the diabatic error and transition probabilities. We visualize the positions of the MZMs during the $X$-gate process by calculating the time-, space-, and energy-resolved local density of states, $LDOS(t,j,\omega)$ Pandey2023_PRB ; Thomale2013 , which should be accessible in time-dependent differential conductance measurements using scanning tunneling spectroscopy Houselt2010 .
@@ Results
Here, we introduce a minimal and experimentally feasible setup consisting of seven interacting quantum dots arranged in a $Y$-shaped geometry (see Fig. 1), suitable for implementing MZM braiding operations and readout. In this configuration, the quantum dots are connected via electrostatically tunable gates, which enable dynamic control over the hopping amplitude $t^{h}(t)$, pairing strength $\Delta(t)$, and on-site chemical potential $\mu_{j}(t)$ Dvir ; Bordin ; Liu2025 ; Pandey2024_PRR , each carrying a time dependence. This high degree of tunability allows for the adiabatic transport of MZMs from one quantum dot to another tenHaaf ; Pandey2025_PRB , thereby enabling both braiding and fusion operations of MZMs. The Hamiltonian for the $Y$-shaped quantum dot array (with $\phi_{I}=0$, $\phi_{II}=\pi/6$, and $\phi_{III}=-\pi/6$ for each leg Harper ) can be divided into four distinct parts. The Hamiltonian for each arm is given by Sekania2017 :
@@
Here, l=(L1)/3l=(L-1)/3, where $L$ is the total number of sites, $c_{j}$ is the fermionic annihilation operator, and $n_{j}=c^{\dagger}_{j}c_{j}$ is the number operator. The parameters $t^{h}_{ij}$ and $\Delta_{ij}(t)$ denote the bond-dependent hopping and pairing amplitudes, respectively. We assume $t^{h}=\Delta$ for our numerical calculations (unless stated otherwise). We employ a smooth time-dependent sine-squared ramp protocol, which has been shown to outperform the linear ramp protocol Pandey2025_PRB ; Sekania2017 . The protocol is defined as
@@
Figure 3: Simulation of Majorana braiding and qubit dynamics in a $Y$-shaped quantum-dot array. a, Initial configuration at $t=0$ with two pairs of spatially separated Majorana zero modes (MZMs), $(\gamma_{a},\gamma_{b})$ and $(\gamma_{c},\gamma_{d})$. b–g, Space- and time-resolved local density of states $LDOS(\omega,t,j)$ at different stages of the braiding sequence. b, At $t=0$, four zero-energy peaks are observed on sites $j=4,5,6,7$, corresponding to the four MZMs. c, At $t=2T/6$, the MZM $\gamma_{b}$ is transferred from site $j=4$ to $j=1$, leading to zero-energy peaks at $j=1,5,6,7$. d, At $t=3T/6$, $\gamma_{c}$ becomes partially delocalized, resulting in finite spectral weight at sites $j=3,4,6$, while sharp peaks remain at $j=1,5,6$. e, At $t=4T/6$, $\gamma_{c}$ becomes localized at site $j=4$, and zero-energy peaks are observed at $j=1,4,5,7$. f, At $t=T$, after completing a single braid, $\gamma_{b}$ and $\gamma_{c}$ are exchanged, as indicated by the restored peaks at $j=4,5,6,7$. g, At $t=2T$, following a full double braid, the system returns to its initial configuration, showing four zero-energy peaks identical to the initial state. h, Time evolution of the diabatic error $f(t)=1-\sum_{n=1}^{4}|\langle\psi(t)|\psi_{n}^{I}\rangle|^{2}$ as a function of scaled time $t/\tau$ for different ramp durations $\tau$. i, Time evolution of the transition probabilities $P_{n}(t)=|\langle\psi_{n}(0)|\psi(t)\rangle|^{2}$ for all four eigenstates ($n=1,2,3,4$), shown for $\tau=50$. j, Final transition probability $P(2T)$ at the end of the protocol ($t=2T$) as a function of switching time $\tau$ for different values of the Coulomb interaction strength $V$.
@@
where $\psi_{n}^{I}$ denotes the instantaneous many-body eigenstates of the system. For small values of $\tau$ (corresponding to fast switching), the diabatic error takes a larger value, indicating transitions to higher-energy states above the gap. Interestingly, we find that for $\tau\geq 40$, the diabatic error remains low ($f(t)\sim 10^{-4}$), showing negligible excitation above the ground-state manifold. For larger $\tau$, the spikes in the diabatic error appear at times when the on-site potentials at the three initial sites $(j=1,2,3)$ are changed simultaneously. The diabatic error for smaller values of switching time leads to larger braiding error and gives a lower bound to the switching time $\tau\gtrsim\hbar/\Delta_{gap}$ Knapp2016 ; Peeters2024 (here $\Delta_{gap}$ is the excitation gap in the topological phase).
@@
Next, we perform a double braiding operation $U^{2}_{bc}$ between MZMs $\gamma_{b}$ and $\gamma_{c}$ (see schematic of Fig. 4(b)). As discussed earlier, this operation transforms the MZM qubit state from $|11\rangle$ to $|00\rangle$. To read out the final state after double braiding (at $t=2T$), we repeat the same $Z$-fusions $F^{Z}_{ba}$ and $F^{Z}_{cd}$. The second row of Figs. 4(g-j) presents the results of $Z$-fusion for different fusion switching times $\tau_{f}$, following a braiding protocol with $\tau=50$ (braiding switching time). In this case, the hole component of the local density of states, $LDOS^{h}(\omega,t,j)$, exhibits a peak with height one at $\omega=0$ (Fig. 4(h)), while the electron component, $LDOS^{e}(\omega,t,j)$, shows no peak at site $j=5$ (Fig. 4(g)). This indicates the formation of a vacuum state ($|00\rangle$) after $Z$-fusion, in contrast to the electron formation $|11\rangle$ observed prior to braiding. Figures 4(i,j), show the time-dependent charge density $\langle n_{j}(t)\rangle$ at sites $j=5$ and $j=7$, respectively, for various $\tau_{f}$. In both cases, the charge density approaches zero for $\tau_{f}\gtrsim 40$, confirming the formation of holes at both sites following fusion. These results validate that the double braid operation leads to a state transformation $|11\rangle\rightarrow|00\rangle$ and should be obtained in the quantum-dot experiment using charge sensor Szechenyi2020 ; Gharavi2016 .
@@ Discussion
Method
Local density-of-states
We have calculated the time-dependent local density of states using the exact diagonalization method Pandey2023_PRB . To simulate the time evolution, we propagate the initial wavefunction $|\Psi(0)\rangle$ using the time-dependent Hamiltonian $H(t)$ up to time $t$:
$|\Psi(t)\rangle=\mathcal{T}\exp\left(-i\int_{0}^{t}H(s)\,ds\right)|\Psi(0)\rangle,$ (4)
where $\mathcal{T}$ denotes the time-ordering operator Kells2014 . We then evaluate the double-time Green’s function Kennes2014 $G(t,t^{\prime})$ using the final Hamiltonian $H_{f}=H(t_{f})$ at the final time $t=t_{f}$. For the electronic component, the Green’s function is given by:
$G^{\text{elec}}_{j}(t,t^{\prime})=\langle\Psi(t)|c_{j}^{\dagger}\,e^{iH_{f}t^{\prime}}\,c_{j}\,e^{-iH_{f}t^{\prime}}|\Psi(t)\rangle.$ (5)
The electronic part of the time-dependent local density of states at site $j$ is obtained by performing a Fourier transform of the Green’s function $G^{\text{elec}}_{j}(t,t^{\prime})$ with respect to $t^{\prime}$:
$\text{LDOS}^{\text{e}}_{j}(\omega,t)=\frac{1}{\pi}\,\text{Im}\left[\int_{0}^{T}dt^{\prime}\,e^{i(\omega+i\eta)t^{\prime}}\,iG^{\text{elec}}_{j}(t,t^{\prime})\right],$ (6)
where the integration is carried out up to $T=60$, and the energy broadening parameter is set to $\eta=0.1$ throughout this work.
In an analogous way, the hole component of the LDOS is derived from the corresponding hole Green’s function:
$G^{\text{hole}}_{j}(t,t^{\prime})=\langle\Psi(t)|c_{j}(t^{\prime})\,c_{j}^{\dagger}|\Psi(t)\rangle,$ (7)
The total local density of states at site $j$ is given by the sum of the electron and hole components:
$\text{LDOS}_{j}(\omega,t)=\text{LDOS}^{\text{e}}_{j}(\omega,t)+\text{LDOS}^{\text{h}}_{j}(\omega,t).$ (8)
Data availability
The data that support the findings of this study are available from the corresponding author upon request.
Code availability
The computer codes used in this study are available from the corresponding author upon request.
Funding Declaration
@@
Author contributions
B.P. designed the project and carried out the numerical calculations. B.P. and E.D. wrote the manuscript. S.O., P.L., and S.L. provided valuable comments and discussions. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.
Additional information
Correspondence and requests for materials should be addressed to Bradraj Pandey (bradraj.pandey@gmail.com).